Made up element Zh

Molar mass= 839.92

85.7 g of Zh2

find the number of moles Zh in compound.


1
Expert's answer
2021-04-19T04:35:41-0400

Molar mass of Zh=839.92

Mass=85.7/2=42.85

Moles=mass/molar mass

42.85/839.92=0.05

Ans=0.05moles